# Teaching Large Language Models to Reason through Learning and Forgetting

### v2

@@ Abstract
Leveraging inference-time search in large language models has proven effective in further enhancing a trained model’s capability to solve complex mathematical and reasoning problems. However, this approach significantly increases computational costs and inference time, as the model must generate and evaluate multiple candidate solutions to identify a viable reasoning path. To address this, we propose an effective approach that integrates search capabilities directly into the model by fine-tuning it using both successful (learning) and failed reasoning paths (forgetting) derived from diverse search methods. While fine-tuning the model with these data might seem straightforward, we identify a critical issue: the model’s search capability tends to degrade rapidly if fine-tuning is performed naively. We show that this degradation can be substantially mitigated by employing a smaller learning rate. Extensive experiments on the challenging Game-of-24 and Countdown mathematical reasoning benchmarks show that our approach not only outperforms both standard fine-tuning and inference-time search baselines but also significantly reduces inference time by 180.111Code is available at https://github.com/twni2016/llm-reasoning-uft.
@@ 1 Introduction
Chain-of-Thought (CoT) and inference-time search (search) have been proposed to enhance LLM generalization to reasoning problems at test-time. Specifically, given a task description and input , CoT (Wei et al., 2022) outputs a single reasoning path via greedy decoding, or samples multiple paths to compute average performance. In contrast, search-based methods (Lightman et al., 2023; Yao et al., 2023a; Snell et al., 2024) explicitly generate intermediate reasoning (possibly with multiple passes) before extracting via a reward model or self-evaluation. Despite inference-time search’s effectiveness, it comes with greater computational costs (e.g., Fig. 1 and Chen et al. (2024b; a)).
A beneficial outcome of utilizing inference-time sampling strategies, such as CoT and search-based methods, is the generation of rich datasets. These datasets can subsequently be fed back into the model for further fine-tuning, potentially enhancing the base model’s overall performance. Prior work has primarily focused on distilling the knowledge from successful reasoning paths, drawing from CoT-generated data (Zelikman et al., 2022; Yuan et al., 2023; Singh et al., 2023), paths from search trees (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a), or full search traces (Gandhi et al., 2024; Lehnert et al., 2024). However, in many complex problems, failed (reasoning) paths far outnumber successful ones, representing an underutilized resource that can be effectively used for training. Additionally, diverse search algorithms yield varying solutions, potentially improving data coverage and reasoning quality. Nonetheless, effectively leveraging both successful and failed reasoning paths and integrating data from both CoT and search methods to enhance reasoning remains largely underexplored.
In this paper, we propose a straightforward and simple pipeline consisting of three phases designed to enhance the reasoning capabilities of large language models: data generation, fine-tuning, and evaluation. First, we collect heterogeneous reasoning data from both CoT and the search-based methods, and convert the data into a unified CoT-style path format. Next, we fine-tune the base LLM by learning successful paths with the supervised fine-tuning (SFT) and forgetting failed paths with the unlikelihood loss. We refer to this approach as unlikelihood fine-tuning (UFT). Finally, we evaluate the fine-tuned model on test sets using both CoT and inference-time search, with a primary focus on efficient CoT inference. We demonstrate the effectiveness of UFT on two math reasoning domains: Game-of-24 and Countdown, using Qwen2.5-Math 1.5B and 7B (Yang et al., 2024). We summarize our key findings below:
@@
Forgetting improves performance: Incorporating failed reasoning paths using UFT improves CoT performance by – on average, with up to gains in best-case scenarios.
@@ 2 Related Work
Inference-time search has shown remarkable success across reasoning (Wang et al., 2022; Stechly et al., 2024; Hao et al., 2024; Snell et al., 2024) and planning tasks (Valmeekam et al., 2023; Zheng et al., 2024; Bohnet et al., 2024). Approaches include best-of-n sampling (Nakano et al., 2021; Wang et al., 2022), multi-agent debates (Du et al., 2023), and iterative correction via predefined rules (Bai et al., 2022) or self-refinement (Shinn et al., 2023; Madaan et al., 2024). A key direction, which this paper focuses on, integrates classic search algorithms with LLMs, employing BFS (Yao et al., 2023a; Xie et al., 2023), DFS (Qin et al., 2023), A* (Zhuang et al., 2023; Meng et al., 2024; Koh et al., 2024) and MCTS (Hao et al., 2023; Zhou et al., 2023; Zhao et al., 2024; Xie et al., 2024; Gao et al., 2024). While effective, inference-time search methods remain computationally expensive: self-refinement requires multiple expensive passes, multi-agent approaches demand substantial GPU memory, and search trees grow exponentially with depth (Chen et al., 2024b).
@@
Synthetic data, generated by classic algorithms or LLMs, is widely used for SFT in reasoning tasks to mitigate the scarcity of human-annotated datasets (Liu et al., 2024). However, despite its abundance, synthetic responses may fail to solve the task. To address this, most prior work distill reasoning knowledge to policy (Hinton et al., 2015) by selecting only correct responses. Policy distillation strategies can be categorized based on whether the source and target data originate from CoT or search algorithms. CoT-to-CoT distillation learns CoT reasoning from correct CoT paths filtered by rewards, widely used in practice (Zelikman et al., 2022; 2024; Uesato et al., 2022; Yuan et al., 2023; Gulcehre et al., 2023; Singh et al., 2023). Search-to-search distillation either learns a meta-policy by imitating entire search traces (Yang et al., 2022; Gandhi et al., 2024; Lehnert et al., 2024) or refines the proposal policy in tree search by imitating extracted paths (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a). Notably, CPO (Zhang et al., 2024c) performs search-to-CoT distillation using preference pairs from Tree-of-Thought without a verifier. In comparison, we focus on CoT, search-to-CoT distillation from diverse reasoners and distill unpaired positive and negative data to the policy, labeled by a rule-based verifier.
@@
The abundance of failed responses compared to successful ones provides an opportunity to leverage them for more efficient and generalizable learning. In case of LLM-based reasoning, a common approach is to train a reward model (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023; Feng et al., 2023; Zhang et al., 2024a; Hosseini et al., 2024) for response re-ranking and search during inference. However, this does not inherently improve the base model as a policy. Another approach extends vanilla DPO (Rafailov et al., 2023) by building preference pairs and fine-tuning the policy with offline preference optimization (Pang et al., 2024; Zhang et al., 2024c; Pal et al., 2024; Setlur et al., 2024). Beyond reasoning, LLM safety literature explores methods for unlearning harmful text (Welleck et al., 2019; Keskar et al., 2019; Yao et al., 2023b; Zhang et al., 2024b). Instead of forming preference pairs or incurring computational overhead to identify high-credit steps, we adopt unlikelihood training (Welleck et al., 2019) to directly forget failed responses as an auxiliary loss, avoiding the limitation of optimizing only relative probabilities within preference pairs (Tuan & Wang, 2024; Feng et al., 2024).
@@ 3 Preliminaries
We formalize a reasoning task as a token-level Markov decision process (MDP) . The initial state consists of a tokenized sequence representing the task description and input. An action is chosen based on the current state , which is the sequence of all previous tokens: . A terminal state is reached upon generating a special end-of-sequence token. The ground-truth reward function is a rule-based verifier that checks the correctness of , yielding , where means failure and means success. The goal is to find a policy that generates an action sequence (denoted as ) which maximizes the reward: . We refer to the terminal state as a (reasoning) path.
@@
CoT directly outputs one (via greedy decoding) or multiple reasoning paths (under non-zero temperature) without intermediate search process222CoT’s performance is often measured by the average success rate over generated reasoning paths, also known as the pass@1 metric (Chen et al., 2021; Guo et al., 2025), as used in this work.. On the contrary, ToT and RAP are inference-time search methods, first constructing a search tree of reasoning paths before selecting a final path as the answer. Their performance is evaluated based on the success rate of the selected reasoning path. Please see Appendix A for more details.
4 Method: Unlikelihood Fine-Tuning for Reasoning
In this section, we describe our approach to solving a reasoning task, as outlined in Fig. 2. Our approach involves three stages: (1) generating reasoning data, (2) fine-tuning the base LLM from unpaired correct and failed paths and (3) evaluating it with reasoning algorithms.
4.1 Data Generation Using LLM and Classic Reasoners
@@
For search-based methods, we extract all root-to-leaf paths from the search tree – rather than only the final selected path – to follow the format of CoT-style paths. This enables (1) unifying heterogeneous search traces into a common format to improve coverage and data quality, and (2) efficient inference, as CoT inference is fastest. We aggregate CoT-style paths from all reasoners to compose the training dataset :
where is the set of considered reasoners, and is the set of all reasoning paths produced by reasoner . Note that classic reasoners do not depend on . Each path is labeled by a verifier as correct or failed , splitting into a correct dataset and a failed dataset .
4.2 Unlikelihood Fine-Tuning from Unpaired Correct and Failed Paths
While prior work typically fine-tune LLMs on correct CoT-generated paths (Yuan et al., 2023; Singh et al., 2023) or on preference pairs of correct and failed responses (Pang et al., 2024; Zhang et al., 2024c), we extend these approaches in two directions: (1) incorporating diverse reasoners beyond CoT to augment the training data, (2) fine-tuning the model to avoid and forget failed reasoning patterns using only failed examples. Unlike preference-based methods, our approach scales better by leveraging unpaired negative data, without requiring success-failure pairs for the same question.
@@ 4.2 Unlikelihood Fine-Tuning from Unpaired Correct and Failed Paths
First, we fine-tune the model using negative log-likelihood (NLL) loss – also known as supervised fine-tuning (SFT) (Ouyang et al., 2022; Cen et al., 2025) – on correct reasoning paths collected from multiple reasoners:
Examining the gradient of the NLL objective (Eq. 2), , reveals its connection to REINFORCE (Williams, 1992) with a binary (indicator) reward function, as noted in prior work (Zelikman et al., 2022; Gulcehre et al., 2023; Singh et al., 2023)333This resemblance holds when is close to on-policy data (Fakoor et al., 2020), e.g., CoT-generated, but may break down as model weights shift during fine-tuning or if includes search-generated data..
@@ 5 Experiments
| Training method | Inference method | Countdown test set | Game-of-24 test set | ||
| Succ rate | Time | Succ rate | Time | ||
| Qwen2.5-Math 7B (base model, no fine-tuning) | CoT (greedy) | 6.2% | 1.0 | 6% | 0.1 |
| CoT (pass@1) | 5.6% | 3.4 | 6% | 0.6 | |
| Search: ToT | 25% | 90 | 28% | 6.5 | |
| Search: RAP | 10.2% | 278 | 27% | 8.3 | |
| SFT baseline (, CoT data only) | CoT (greedy) | 32.5% | 0.5 | 16% | 0.1 |
| CoT (pass@1) | 32.2% | 3.1 | 12.3% | 0.6 | |
| Search: ToT | 22.2% | 96 | 30.7% | 8 | |
| Search: RAP | 17.5% | 141 | 27.7% | 16 | |
| Ours (, CoT or All data) | CoT (greedy) | 57.4% | 0.5 | 42% | 0.1 |
| CoT (pass@1) | 53.5% | 3.4 | 29.7% | 0.8 | |
| Search: ToT | 28.6% | 153 | 24% | 13 | |
| Search: RAP | 28.9% | 439 | 33.7% | 75 | |
@@ 5.1 Setup
We mainly build upon the LLM-reasoners repository666https://github.com/maitrix-org/llm-reasoners (Hao et al., 2024), which provides a unified library for all the LLM reasoners studied here. We significantly accelerate inference by batching inputs using vLLM (Kwon et al., 2023). All reported inference time in our results is based on our optimized codebase. For each LLM reasoner algorithm, we vary its hyperparameters to gather data from . For CoT (Wei et al., 2022), we vary the temperature between and top-p between to have 9 variants. For each variant, we use a fixed 5-shot prompt template and collect paths for each case in Game-of-24 and paths for Countdown. For Game-of-24, we follow LLM-reasoners to implement ToT (Yao et al., 2023a) (varying the beam size from to ) and RAP (Hao et al., 2023) (varying the exploration parameter from to ). Since Countdown is implemented in Stream-of-Search (SoS) (Gandhi et al., 2024), we use SoS code to implement classic BFS and DFS algorithms, which do not use LLMs but rely on the oracle verifier. We merge the reasoning paths generated by all the variants of the same reasoner, and then remove duplicates by exact matching, to serve as our training dataset. See Appendix A for all details.
Step 2: Fine-tuning setup.
We follow the alignment handbook (Tunstall et al., 2024) to implement standard SFT and UFT on an instance with 8 A100 GPUs. We fix the batch size to and train epochs on each dataset for Game-of-24 and epochs for Countdown. We use a cosine schedule and sweep the peak learning rate over (1e-5, 5e-6, 2e-6) for Q1.5B and over (5e-6, 2e-6, 1e-6) for Q7B. We set to have standard SFT (NLL) and sweep over (1e-3, 1e-4, 1e-5, 1e-6) for UFT. See Appendix B for details.
@@
| Base | Countdown training data | Countdown test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 20.0% | 2.2% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.3% | ||||||
| Qwen2.5- Math 7B | CoT | 37.6% | 25% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.9% | ||||||
| Base | Countdown training data | Game-of-24 test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 20.0% | 8% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.3% | ||||||
| Qwen2.5- Math 7B | CoT | 37.6% | 28% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.9% | ||||||
| Base | Game-of-24 training data | Game-of-24 test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 82.1% | 8% | ||||
| ToT+RAP | 44.7% | ||||||
| CoT+ToT+RAP | 87.6% | ||||||
| Qwen2.5- Math 7B | CoT | 92.9% | 28% | ||||
| ToT+RAP | 69.3% | ||||||
| CoT+ToT+RAP | 95.3% | ||||||

### v3

# Offline Learning and Forgetting for Reasoning with Large Language Models

@@ Abstract
Leveraging inference-time search in large language models has proven effective in further enhancing a trained model’s capability to solve complex mathematical and reasoning problems. However, this approach significantly increases computational costs and inference time, as the model must generate and evaluate multiple candidate solutions to identify a viable reasoning path. To address this, we propose an effective approach that integrates search capabilities directly into the model by fine-tuning it on unpaired successful (learning) and failed reasoning paths (forgetting) derived from diverse search methods. A key challenge we identify is that naive fine-tuning can degrade the model’s search capability; we show this can be mitigated with a smaller learning rate. Extensive experiments on the challenging Game-of-24 and Countdown reasoning benchmarks show that, replacing CoT-generated data with search-generated data for offline fine-tuning improves success rates by around 23% over inference-time search baselines, while reducing inference time by 180. On top of this, our learning and forgetting objective consistently outperforms both supervised fine-tuning and preference-based methods.111Code is available at https://github.com/twni2016/llm-reasoning-uft.
@@ 1 Introduction
Chain-of-Thought (CoT) and inference-time search (search) have been proposed to enhance LLM generalization to reasoning problems at test time. Specifically, given a task description and input , CoT (Wei et al., 2022) outputs a single reasoning path via greedy decoding, or samples multiple paths to compute average performance. In contrast, search-based methods (Lightman et al., 2023; Yao et al., 2023a; Snell et al., 2024) explicitly generate intermediate reasoning (possibly with multiple passes) before extracting via a reward model or self-evaluation. Despite inference-time search’s effectiveness, it comes with greater computational costs (e.g., Fig. 1 and Chen et al. (2024b; a)).
A beneficial outcome of utilizing inference-time sampling strategies, such as CoT and search-based methods, is the generation of rich datasets. These datasets can subsequently be fed back into the model for further fine-tuning, potentially enhancing the base model’s overall performance. Prior work has primarily focused on distilling the knowledge from successful reasoning paths, drawing from CoT-generated data (Zelikman et al., 2022; Yuan et al., 2023b; Singh et al., 2023), paths from search trees (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a), or full search traces (Gandhi et al., 2024; Lehnert et al., 2024). However, in many complex problems, failed (reasoning) paths far outnumber successful ones, representing an underutilized resource. While potentially valuable, these failed paths are often unpaired, making them difficult to use in preference optimization (Rafailov et al., 2023). Diverse search algorithms further enrich data coverage and reasoning quality but produce heterogeneous structures (e.g., trees) unlike linear CoT traces. This raises a natural question: can we make full use of both successful and failed reasoning paths, from both CoT and search-based methods, to improve LLM reasoning?
In this paper, we propose a straightforward method consisting of three phases designed to enhance the reasoning capabilities of large language models: data generation, fine-tuning, and evaluation. First, we collect heterogeneous reasoning data from both CoT and the search-based methods, and convert the data into a unified CoT-style path format. Next, we fine-tune the base LLM by learning successful paths with the supervised fine-tuning (SFT) and forgetting unpaired failed paths with the unlikelihood loss. We refer to this approach as unlikelihood fine-tuning (UFT). Finally, we evaluate the fine-tuned model on test sets using both CoT and inference-time search, with a primary focus on efficient CoT inference. We demonstrate the effectiveness of UFT on two math reasoning domains: Game-of-24 and Countdown, using Qwen2.5-Math 1.5B and 7B (Yang et al., 2024). We summarize our key findings below:
@@
Unpaired forgetting reliably improves performance: Incorporating failed paths using UFT improves CoT performance by – on average, with up to gains in the best case. In contrast, preference-based methods like CPO-SimPO (Xu, 2024) often underperform SFT due to the paired format.
@@ 2 Related Work
Inference-time search has shown remarkable success across reasoning (Wang et al., 2022; Stechly et al., 2024; Hao et al., 2024; Snell et al., 2024) and planning tasks (Valmeekam et al., 2023; Zheng et al., 2024; Bohnet et al., 2024). Approaches include best-of-n sampling (Nakano et al., 2021; Wang et al., 2022), multi-agent debates (Du et al., 2023), and iterative correction via predefined rules (Bai et al., 2022) or self-refinement (Shinn et al., 2023; Madaan et al., 2024). A key direction, which this paper focuses on, integrates classic search algorithms with LLMs, employing BFS (Yao et al., 2023a; Xie et al., 2023), DFS (Qin et al., 2023), A* (Zhuang et al., 2023; Meng et al., 2024a; Koh et al., 2024) and MCTS (Hao et al., 2023; Zhou et al., 2023; Zhao et al., 2024; Xie et al., 2024; Gao et al., 2024). While effective, inference-time search methods remain computationally expensive: self-refinement requires multiple expensive passes, multi-agent approaches demand substantial GPU memory, and search trees grow exponentially with depth (Chen et al., 2024b).
@@
Synthetic data, generated by classic algorithms or LLMs, is widely used for SFT in reasoning tasks to mitigate the scarcity of human-annotated datasets (Liu et al., 2024). However, despite its abundance, synthetic responses may fail to solve the task. To address this, most prior work distill reasoning knowledge to policy (Hinton et al., 2015) by selecting only correct responses. Policy distillation strategies can be categorized based on whether the source and target data originate from CoT or search algorithms. CoT-to-CoT distillation learns CoT reasoning from correct CoT paths filtered by rewards, widely used in practice (Zelikman et al., 2022; 2024; Uesato et al., 2022; Yuan et al., 2023b; Gulcehre et al., 2023; Singh et al., 2023). Search-to-search distillation either learns a meta-policy by imitating entire search traces (Yang et al., 2022; Gandhi et al., 2024; Lehnert et al., 2024) or refines the proposal policy in tree search by imitating extracted paths (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a). Notably, Zhang et al. (2024c) performs search-to-CoT distillation using preference pairs from Tree-of-Thought without a verifier. In comparison, we focus on CoT, search-to-CoT distillation from diverse reasoners and distill unpaired positive and negative data to the policy, labeled by a rule-based verifier.
@@
Negative data has been extensively studied in LLM safety, where the goal is to unlearn harmful content using objectives such as gradient ascent and unlikelihood training (Welleck et al., 2019; Keskar et al., 2019; Yao et al., 2023b; Zhang et al., 2024b). In preference optimization, negative examples are used to increase the relative likelihood of preferred responses in paired data , commonly used in alignment (Rafailov et al., 2023; Yuan et al., 2023a; Zhao et al., 2023; Hong et al., 2024; Xu et al., 2024; Meng et al., 2024b) and reasoning (Pal et al., 2024; Pang et al., 2024; Zhang et al., 2024c; Setlur et al., 2024). However, this paired format discards unpaired positives or negatives, reducing data efficiency. KTO (Ethayarajh et al., 2024) addresses the limitation of relative likelihood (Tuan & Wang, 2024) by learning from unpaired data, although it requires a reference model. In LLM reasoning, negative data is also used to train a reward model (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023; Feng et al., 2023; Zhang et al., 2024a; Hosseini et al., 2024) for response re-ranking and search during inference. However, this does not inherently improve the base model as a policy. Instead of requiring preference pairs or reference models, we adopt unlikelihood training (Welleck et al., 2019) to directly forget failed paths as an auxiliary loss.
@@ 3 Preliminaries
We formalize a reasoning task as a token-level Markov decision process (MDP) (Sutton et al., 1998) . The initial state consists of a tokenized sequence representing the task description and input. An action is chosen based on the current state , which is the sequence of all previous tokens: . A terminal state is reached upon generating a special end-of-sequence token. The ground-truth reward function is a rule-based verifier that checks the correctness of , yielding , where means failure and means success. The goal is to find a policy that generates an action sequence (denoted as ) which maximizes the reward: . We refer to the terminal state as a (reasoning) path.
@@
CoT directly outputs one (via greedy decoding) or multiple reasoning paths (under non-zero temperature) without intermediate search process222CoT’s performance is often measured by the average success rate over generated reasoning paths, also known as the pass@1 metric (Chen et al., 2021; Guo et al., 2025), as used in this work.. On the contrary, ToT and RAP are inference-time search methods, first constructing a search tree of reasoning paths before selecting a final path as the answer. Their performance is evaluated based on the success rate of the selected reasoning path. Please see Appendix C for more details.
4 Fine-tuning on Unpaired Correct and Failed Paths from Diverse Reasoners
In this section, we describe our approach to solving a reasoning task, as outlined in Fig. 2. Our approach involves three stages: (1) generating reasoning data from diverse reasoners, (2) offline fine-tuning the base LLM from unpaired correct and failed paths, and (3) evaluating it with reasoning algorithms.
4.1 CoT-style Data Generation Using LLM and Classic Reasoners
@@
For search-based methods, we extract all root-to-leaf paths from the search tree – rather than only the final selected path – to follow the format of CoT-style paths. This enables (1) unifying heterogeneous search traces into a common format as training data, and (2) efficient inference, as CoT inference is fastest. We aggregate CoT-style paths from all reasoners to compose the training dataset :
where is the set of considered reasoners, and is the set of all reasoning paths produced by reasoner . Note that classic reasoners do not depend on . Each path is labeled by a verifier as correct or failed , splitting into two unpaired datasets: the correct dataset and the failed dataset .
4.2 Unlikelihood Fine-Tuning on Unpaired Correct and Failed Paths
While prior work typically fine-tune LLMs on correct CoT-generated paths (Yuan et al., 2023b; Singh et al., 2023) or on preference pairs of correct and failed responses (Pang et al., 2024; Zhang et al., 2024c), we extend these approaches in two directions: (1) incorporating diverse reasoners beyond CoT to augment the training data, (2) fine-tuning the model to avoid and forget failed reasoning patterns using only failed examples. Unlike most preference-based methods, our approach leverages unpaired positive or negative data, without requiring success-failure pairs for the same question.
@@ 4.2 Unlikelihood Fine-Tuning on Unpaired Correct and Failed Paths
First, we consider the negative log-likelihood (NLL) loss, also known as supervised fine-tuning (SFT) (Ouyang et al., 2022; Cen et al., 2025), on correct reasoning paths collected from multiple reasoners:
Examining the gradient of the NLL objective (Eq. 2), , reveals its connection to REINFORCE (Williams, 1992) with a binary (indicator) reward function, as noted in prior work (Zelikman et al., 2022; Gulcehre et al., 2023; Singh et al., 2023)333This resemblance holds when is close to on-policy data (Fakoor et al., 2020), e.g., CoT-generated, but may break down as model weights shift during fine-tuning or if includes search-derived data..
@@ 5 Experiments
Test set & Inference | Base LLM | lr=5e-6, CoT+BFS+DFS training data | lr=1e-6, CoT training data | ||||
SFT (=0) | CPO-SimPO | UFT (=1e-3) | SFT (=0) | CPO-SimPO | UFT (=1e-4) | ||
Countdown (1000 cases) | |||||||
CoT (greedy) | 6.2% / 1.0m | 55.4% / 0.6m | 50.6% / 0.7m | 57.4% / 0.5m | 24.7% / 0.5m | 23.4% / 0.8m | 24.7% / 0.5m |
CoT (pass@1) | 5.6% / 3.4m | 51.8% / 3.5m | 45.7% / 4.7m | 53.5% / 3.4m | 24.1% / 2.9m | 22.1% / 5.7m | 24.7% / 3.3m |
search: ToT | 25.0% / 90m | 2.4% / 104m | 4.1% / 61m | 0.0% / 1.9m | 22.2% / 96m | 23.3% / 179m | 28.6% / 153m |
search: RAP | 10.2% / 278m | 7.3% / 155m | 7.0% / 137m | 0.0% / 341m | 17.5% / 141m | 17.7% / 190m | 28.9% / 439m |
Game-of-24 (100 cases) | |||||||
CoT (greedy) | 6.0% / 0.1m | 35.0% / 0.1m | 24.7% / 0.1m | 42.0% / 0.1m | 8.0% / 0.1m | 6.7% / 0.1m | 8.3% / 0.1m |
CoT (pass@1) | 6.0% / 0.6m | 28.3% / 0.7m | 23.0% / 1.3m | 29.7% / 0.8m | 7.3% / 0.6m | 7.0% / 1.5m | 6.7% / 0.8m |
search: ToT | 28.0% / 6.5m | 2.3% / 8.2m | 1.0% / 4.3m | 0.3% / 0.4m | 30.7% / 8m | 22.7% / 13m | 24.0% / 13m |
search: RAP | 27.0% / 8.3m | 1.3% / 13m | 3.0% / 19m | 3.7% / 29m | 27.7% / 16m | 25.7% / 42m | 33.7% / 75m |
@@ 5.1 Setup
We mainly build upon the LLM-reasoners repository666https://github.com/maitrix-org/llm-reasoners (Hao et al., 2024), which provides a unified library for all the LLM reasoners studied here. We significantly accelerate inference by batching inputs using vLLM (Kwon et al., 2023). All reported inference time in our results is based on our optimized codebase. For each LLM reasoner algorithm, we vary its hyperparameters to gather data from . For CoT (Wei et al., 2022), we vary the temperature between and top-p between to have 9 variants. For each variant, we use a fixed 5-shot prompt template and collect paths for each case in Game-of-24 and paths for Countdown. For Game-of-24, we follow LLM-reasoners to implement ToT (Yao et al., 2023a) (varying the beam size from to ) and RAP (Hao et al., 2023) (varying the exploration parameter from to ). Since Countdown is implemented in Stream-of-Search (SoS) (Gandhi et al., 2024), we use SoS code to implement classic BFS and DFS algorithms, which do not use LLMs but rely on the oracle verifier. We merge the reasoning paths generated by all the variants of the same reasoner, and then remove duplicates by exact matching, to serve as our training dataset. See Appendix C for all details.
Step 2: Fine-tuning setup and baselines.
We follow the alignment handbook (Tunstall et al., 2024) to implement all fine-tuning methods on an instance with 8 A100 GPUs. All methods share the same datasets and use common hyperparameters when applicable. We fix the batch size to and train epochs on each dataset for Game-of-24 and epochs for Countdown. We use a cosine schedule and sweep the peak learning rate over (1e-5, 5e-6, 2e-6) for Q1.5B and over (5e-6, 2e-6, 1e-6) for Q7B. We set to have the SFT baseline and sweep over (1e-3, 1e-4, 1e-5, 1e-6) for our UFT. We also implement two preference-based baselines: SimPO (Meng et al., 2024b), which optimizes over paired preference data constructed from successful and failed reasoning paths, and CPO-SimPO (Xu, 2024), which augments SimPO with NLL term to better support reasoning tasks. Both methods are reference-free and leverage negative data as UFT. Notably, CPO-SimPO closely resembles the non-iterative variant of RPO (Pang et al., 2024). See Appendix A for details and Table 4 for method comparison.
@@
SFT | SimPO | CPO-SimPO | UFT | |
CoT data (quality: 37.6%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) | ||||
BFS+DFS data (quality: 86.2%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) | ||||
CoT+BFS+DFS data (quality: 86.9%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) |
SFT | SimPO | CPO-SimPO | UFT | |
CoT data (quality: 20.0%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) | ||||
BFS+DFS data (quality: 86.2%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) | ||||
CoT+BFS+DFS data (quality: 86.3%) | ||||
Countdown succ (greedy) | ||||
Countdown succ (pass@1) | ||||
Game-of-24 succ (greedy) | ||||
Game-of-24 succ (pass@1) |
SFT | SimPO | CPO-SimPO | UFT | |
CoT data (quality: 92.9%) | ||||
succ (greedy) | ||||
succ (pass@1) | ||||
ToT+RAP data (quality: 69.3%) | ||||
succ (greedy) | ||||
succ (pass@1) | ||||
CoT+ToT+RAP data (quality: 95.3%) | ||||
succ (greedy) | ||||
succ (pass@1) |
SFT | SimPO | CPO-SimPO | UFT | |
CoT data (quality: 82.1%) | ||||
succ (greedy) | ||||
succ (pass@1) | ||||
ToT+RAP data (quality: 44.7%) | ||||
succ (greedy) | ||||
succ (pass@1) | ||||
CoT+ToT+RAP data (quality: 87.6%) | ||||
succ (greedy) | ||||
succ (pass@1) |